# thanks: Corresponding author

### v1

@@ I Introduction
Originating from the hyperfine structure transition of neutral hydrogen atoms, the 21-cm signal constitutes a pivotal tool for investigating the first billion years of the universe including the stages of cosmic dark ages, cosmic dawn, and the epoch of reionization (EoR). During the phase when reionization is still incomplete, a significant amount of neutral hydrogen atoms accumulate within and around small-scale structures such as mini dark halos and dwarf galaxies. If the light from high-redshift radio-bright sources, such as radio-loud quasars and the radio afterglows of gamma-ray bursts, happens to traverse these structures along the line of sight, the 21-cm absorption of neutral hydrogen will produce narrow and dense absorption lines in their spectra. This phenomenon is known as the “21-cm forest” Carilli et al. (2002); Furlanetto and Loeb (2002); Furlanetto (2006a); Xu et al. (2009, 2010, 2011); Ciardi et al. (2013). The absorption line characteristics of the 21-cm forest can elucidate the small-scale distribution of neutral hydrogen throughout the EoR and yield vital insights into the heating processes and the formation of small-scale structures in the early universe.
@@
Dark matter (DM) is instrumental in the formation of cosmic structures. In contrast to the cold dark matter (CDM) model, non-CDM models, including warm dark matter (WDM) and ultra-light dark matter, exhibit distinct behaviors on small scales Bode et al. (2001); Marsh (2016); Muñoz et al. (2020); Jones et al. (2021); Ferreira (2021); Shen et al. (2023). Specifically, in the WDM model, the free-streaming of dark matter particles inhibits the formation of small-scale structures, resulting in a scarcity of dark matter halos Avila-Reese et al. (2001); Smith and Markovic (2011); Schneider et al. (2013). Since the 21-cm forest signal is predominantly sourced by these low-mass dark matter halos, different dark matter models directly impact the abundance of 21-cm forest absorption lines Sitwell et al. (2014); Shimabukuro et al. (2014). Furthermore, early X-ray sources in the universe heat the intergalactic medium (IGM), subsequently reducing the intensity and number of 21-cm signals Xu et al. (2009); Mack and Wyithe (2012); Shimabukuro et al. (2014). By measuring the one-dimensional (1D) power spectrum along the line of sight, it is feasible to disentangle these two effects, as their impacts on various scales are discernible. Consequently, the 21-cm forest serves as a dual probe for exploring the properties of dark matter and the heating levels in the early universe Shao et al. (2023).
@@
Despite the theoretical promise of the 21-cm forest, there remains a scarcity of analytical models for this probe. Direct modeling of the 21-cm forest is intricate due to the interplay of multiple physical processes, including dark matter halo formation, gas thermal evolution, and the effects of the radiation background Furlanetto and Loeb (2002); Xu et al. (2011); Šoltinský et al. (2021, 2022). Conversely, the 1D power spectrum of the 21-cm forest encapsulates extensive information about the matter power spectrum, and the analytical modeling of matter power spectra is relatively mature Ma and Fry (2000); Smith et al. (2003); Mead et al. (2015); Philcox et al. (2020); Acuto et al. (2021). Therefore, the analytical modeling of the 1D power spectrum presents a viable approach.
@@
Radio-loud quasars are the most suitable background sources for detecting the 21-cm forest. Numerous studies have already predicted the number of radio-loud quasars observable at high redshifts Haiman et al. (2004); Gloudemans et al. (2021). Recently, Niu et al. Niu et al. (2024) updated the distribution of radio-loud quasars. Utilizing this updated methodology, we simulate the high-redshift radio-loud quasars observable by the SKA. We employ the halo model methodology, coupled with simulated high-redshift radio-loud quasars as background sources, to compute the 1D power spectrum of the 21-cm forest along the line of sight. We specifically consider the ramifications of different DM models (encompassing CDM and WDM) and varying heating levels (primarily X-ray heating). By juxtaposing our results with simulated data, we validate the reliability of our analytical model. This model facilitates an understanding of the perturbations in the 21-cm signal at small scales, thereby aiding in the exploration of dark matter properties and heating processes in the early universe.
@@ II Halo model formalism
The 21-cm photons emmitted by high-redshift background sources are absorbed by neutral hydrogen in and around mini halos (or dwarf galaxies), generating the 21-cm forest signal. The differential brightness temperature observed for the 21-cm signal is commonly expressed as
$\delta T_{\mathrm{b}}(z)$ $\ \approx\frac{T_{\mathrm{S}}(z)-T_{\gamma}(z)}{1+z}\tau(z),$ (1)
where $T_{\mathrm{S}}$ denotes the spin temperature of the HI gas and $\tau$ represents the 21-cm optical depth. $T_{\gamma}$ is the brightness temperature of the background radiation, including the background point source temperature $T_{\mathrm{rad}}$ and the cosmic microwave background (CMB) temperature $T_{\mathrm{CMB}}$. The 21-cm optical depth, which quantifies the absorption strength, can be expressed in terms of the average gas properties as follows Furlanetto (2006a); Madau et al. (1997),
$\tau(z)$ $\approx 0.0085[1+\delta(z)](1+z)^{3/2}\left[\frac{x_{\mathrm{HI}}(z)}{T_{\mathrm{S}}(z)}\right]$ (2)
×$\times\left[\frac{H(z)/(1+z)}{\mathrm{d}v_{\|}/\mathrm{d}r_{\|}}\right]\left(\frac{\Omega_{\rm b}h^{2}}{0.022}\right)\left(\frac{0.14}{\Omega_{\rm m}h^{2}}\right).$
Here, $\delta(z)$ is the gas overdensity, $x_{\mathrm{HI}}(z)$ is the fraction of neutral hydrogen, $H(z)$ is the Hubble parameter, and $\mathrm{d}v_{\|}/\mathrm{d}r_{\|}$ is the velocity gradient along the line of sight. For the cosmological parameters, we set the values as follows: $h$ = 0.6736, $\Omega_{\rm m}$ = 0.3153 and $\Omega_{\rm b}h^{2}$ = 0.02236.
The observed brightness temperature of a background source, at a frequency $\nu$, is related to the flux density $S_{\mathrm{rad}}(\nu)$ by
$T_{\mathrm{rad}}(\nu,z=0)=\frac{c^{2}}{2k_{\rm B}\nu^{2}}\frac{S_{\mathrm{rad}}(\nu)}{\Omega}.$ (3)
Here, $c$ represents the speed of light, $k_{\rm B}$ is the Boltzmann constant, and $\Omega$ denotes the solid angle subtended by the telescope. The flux density $S_{\mathrm{rad}}(\nu)$ of the background source is modeled as a power-law spectrum with a spectral index $\eta$. This model is scaled to 150 MHz, represented as $S_{\text{rad }}(\nu)=S_{150}\left(\frac{\nu}{\nu_{150}}\right)^{\eta}$ Thyagarajan (2020), where $S_{150}$ represents the flux density at 150 MHz, with $\eta=-1.05$ Carilli et al. (2002).
@@
$\delta T_{\mathrm{b}}(z)\approx T_{0}(z)\frac{1+\delta(z)}{T_{\rm K}(z)},$ (4)
where $T_{0}(z)=-0.0085(1+z)^{1/2}T_{\gamma}(z)$ is only related to the brightness temperature of the background radiation.
@@
$P(k_{\|},z)=T_{0}^{2}(z)P_{21}(k_{\|},z),$ (5)
where $P_{21}(k_{\|},z)$ is the 1D power spectrum of the term $[1+\delta(z)]/T_{\rm K}(z)$, related to averaged three-dimensional (3D) power specturm of 21-cm signal $P_{21}(k,z)$,
$P_{21}(k_{\|},z)=\frac{1}{2\pi}\int_{k}^{\infty}kP_{21}(k,z){\rm d}k.$ (6)
@@ II.2 Halo model
where $P_{21}^{1h}(k,z)$ represents the contribution from a single halo and $P_{21}^{2h}(k,z)$ captures the contribution from two separate halos. The specific expressions for each term are given as follows,
@@
where $\rho_{21}(r,z,M)$ is the profile of $[1+\delta(z)]/T_{\rm K}(z)$, including both the density profile and the temperature profile. In the following sections, we will expand on these components, detailing the specific forms of the density and temperature profiles.
@@ II.2.1 Halo mass function
Figure 1 illustrates the halo mass functions for CDM and WDM models at different redshifts. Since the 21-cm forest signal is mainly contributed by low-mass halos, we focus on halos with a minimum mass of $M_{\rm min}=10^{5}M_{\odot}$, corresponding to the Jeans mass scale for the redshifts of interest. The upper mass limit, $M_{4}$, is associated with a virial temperature of $T_{\rm vir}=10^{4}$ K. Within this mass range, the WDM model predicts fewer halos than the CDM model.
Figure 1: Halo mass function in CDM model (solid lines) and WDM model (dashed lines) with $m_{\rm W}$ = 6 keV at different redshifts. The blue and green lines correspond to $z$ = 6 and 10, respectively.
@@ II.2.2 Density profile
Due to the gravitational influence, the gas density in the surroundings of halos is enhanced. Outside the virial radius, we assume the gas density profile aligns with that of the dark matter and can be determined using the infall model Barkana (2004). The gas density profiles in and around halos of different masses are plotted in the left panel of Fig. 2.
@@ II.2.3 Temperature profile
where $f_{\star}$ is the star formation efficiency, $f_{\rm coll}$ is the fraction of matter collapsed into atomic-cooling halos and $f_{\rm X}$ is a normalization factor representing the uncertain X-ray productivity in the early universe.
@@
The heating due to X-rays is dependent on the parameter $f_{\rm X}$, which characterizes the efficiency of X-ray production from early sources. For higher values of $f_{\rm X}$, the gas temperature in neutral regions is significantly raised, reducing the contrast between dense regions and the IGM in terms of their 21-cm absorption signatures. We show the temperature profiles for an un-heated IGM with different masses in the right panel of Fig. 2. According to the recent Hydrogen Epoch of Reionization Array findings Abdurashidova et al. (2023), the temperature of IGM at redshift $z\sim$ 8 is constrained within the range 15.6 K $<T_{\rm K}<$ 656.7 K, corresponding to $0.60.02<f_{\rm X}<$ 0.6. Thus for subsequent analysis, we adopt $f_{\rm X}$ = 0.1 as the fiducial model.
@@ III Forecasts
III.1 Radio-load quasar population
We employ the distribution of high-redshift quasars in Ref. Niu et al. (2024) to simulate background point sources. The abundance of radio-loud quasars can be calculated as follows,

### v2

@@ I Introduction
Originating from the hyperfine structure transition of neutral hydrogen atoms, the 21-cm signal constitutes a pivotal tool for investigating the first billion years of the Universe including the stages of cosmic dark ages, cosmic dawn, and the epoch of reionization (EoR). During the phase when reionization is still incomplete, a significant amount of neutral hydrogen atoms accumulate within and around small-scale structures such as minihalos and dwarf galaxies Furlanetto and Loeb (2002); Xu et al. (2011); Meiksin (2011). If the light from high-redshift radio-bright sources, such as radio-loud quasars Banados et al. (2021); Ighina et al. (2021); Haiman et al. (2004); Gloudemans et al. (2022) and the radio afterglows of gamma-ray bursts Ioka and Meszaros (2005); Salvaterra et al. (2009); Cucchiara et al. (2011); Ciardi et al. (2015); Kinugawa et al. (2019), happens to traverse these structures along the line of sight, the 21-cm absorption of neutral hydrogen will produce narrow and dense absorption lines in their spectra. This phenomenon is known as the “21-cm forest” Carilli et al. (2002); Furlanetto and Loeb (2002); Furlanetto (2006a); Xu et al. (2009, 2010, 2011); Ciardi et al. (2013). The absorption line characteristics of the 21-cm forest can elucidate the small-scale distribution of neutral hydrogen throughout the EoR and yield vital insights into the heating processes and the formation of small-scale structures in the early Universe.
@@
Dark matter (DM) is instrumental in the formation of cosmic structures. In contrast to the cold dark matter (CDM) model, non-CDM models, including warm dark matter (WDM) and ultralight dark matter, exhibit distinct behaviors on small scales Bode et al. (2001); Marsh (2016); Muñoz et al. (2020); Jones et al. (2021); Ferreira (2021); Shen et al. (2023); Shimabukuro et al. (2020); Kawasaki et al. (2020). Specifically, in the WDM model, the free-streaming of dark matter particles inhibits the formation of small-scale structures, resulting in a scarcity of dark matter halos Avila-Reese et al. (2001); Smith and Markovic (2011); Schneider et al. (2013). Since the 21-cm forest signal is predominantly sourced by these low-mass dark matter halos, different dark matter models directly impact the abundance of 21-cm forest absorption lines Sitwell et al. (2014); Shimabukuro et al. (2014). Furthermore, early x-ray sources in the Universe heat the intergalactic medium (IGM), subsequently reducing the intensity and number of 21-cm signals Xu et al. (2009); Mack and Wyithe (2012); Shimabukuro et al. (2014). By measuring the one-dimensional (1D) power spectrum along the line of sight, it is feasible to disentangle these two effects, as their impacts on various scales are discernible. Consequently, the 21-cm forest serves as a dual probe for exploring the properties of dark matter and the heating levels in the early Universe Shao et al. (2023).
@@
Despite the theoretical promise of the 21-cm forest, there remains a scarcity of analytical models for this probe. Direct modeling of the 21-cm forest is intricate due to the interplay of multiple physical processes, including dark matter halo formation, gas thermal evolution, and the effects of the radiation background Furlanetto and Loeb (2002); Ciardi and Ferrara (2005); Xu et al. (2011); Šoltinský et al. (2022). As upcoming observations become increasingly feasible, the lack of analytical models capable of linking observational data to key physical parameters such as the temperature of the IGM or the mass of WDM particles would make it necessary to rely on computationally expensive small-scale simulations. Therefore, it is crucial to develop simple yet effective analytical frameworks that can bridge observational quantities with underlying physical parameters. Conversely, the 1D power spectrum of the 21-cm forest encapsulates extensive information about the matter power spectrum, and the analytical modeling of matter power spectra is relatively mature Ma and Fry (2000); Smith et al. (2003); Mead et al. (2015); Philcox et al. (2020); Acuto et al. (2021). Therefore, the analytical modeling of the 1D power spectrum presents a viable approach.
@@
Radio-loud quasars are the most suitable background sources for detecting the 21-cm forest. Numerous studies have already predicted the number of radio-loud quasars observable at high redshifts Haiman et al. (2004); Wilman et al. (2008); Gloudemans et al. (2021). Recently, Niu et al. Niu et al. (2024) updated the distribution of radio-loud quasars. Utilizing this updated methodology, we simulate the high-redshift radio-loud quasars observable by the SKA. We employ the halo model methodology, coupled with simulated high-redshift radio-loud quasars as background sources, to compute the 1D power spectrum of the 21-cm forest along the line of sight. We specifically consider the ramifications of different DM models (encompassing CDM and WDM) and varying heating levels (primarily x-ray heating). By juxtaposing our results with simulated data, we validate the reliability of our analytical model. This model facilitates an understanding of the perturbations in the 21-cm signal at small scales, thereby aiding in the exploration of dark matter properties and heating processes in the early Universe.
@@ II Halo model formalism
The 21-cm photons emitted by high-redshift background sources are absorbed by neutral hydrogen in and around minihalos (or dwarf galaxies), generating the 21-cm forest signal. The differential brightness temperature observed for the 21-cm signal is commonly expressed as
$\delta T_{\mathrm{b}}(\hat{\bm{n}},z)$ $\ \approx\frac{T_{\mathrm{S}}(\hat{\bm{n}},z)-T_{\gamma}(\hat{\bm{n}},z)}{1+z}\tau(\hat{\bm{n}},z),$ (1)
where $\hat{\bm{n}}$ represents the specific direction of the background source, $T_{\mathrm{S}}$ denotes the spin temperature of the HI gas and $\tau$ represents the 21-cm optical depth. $T_{\gamma}$ is the brightness temperature of the background radiation, including the background point source temperature $T_{\mathrm{point}}$ and the cosmic microwave background (CMB) temperature $T_{\mathrm{CMB}}$. The 21-cm optical depth, which quantifies the absorption strength, can be expressed in terms of the average gas properties as follows Furlanetto (2006a); Madau et al. (1997):
@@
$\times\left[\frac{H(z)/(1+z)}{\mathrm{d}v_{\|}/\mathrm{d}r_{\|}}\right]\left(\frac{\Omega_{\rm b}h^{2}}{0.022}\right)\left(\frac{0.14}{\Omega_{\rm m}h^{2}}\right).$
Here, $\delta(\hat{\bm{n}},z)$ is the gas overdensity, $x_{\mathrm{HI}}(\hat{\bm{n}},z)$ is the fraction of neutral hydrogen, $H(z)$ is the Hubble parameter, and $\mathrm{d}v_{\|}/\mathrm{d}r_{\|}$ is the velocity gradient along the line of sight. For the cosmological parameters, we set the values as follows: $h=0.6736$, $\Omega_{\rm m}=0.3153$, and $\Omega_{\rm b}h^{2}=0.02236$ Aghanim et al. (2020).
The observed brightness temperature of a background source, at a frequency $\nu$, is related to the flux density $S_{\mathrm{point}}(\nu)$ by
$T_{\mathrm{point}}(\hat{\bm{n}},\nu,z=0)=\frac{c^{2}}{2k_{\rm B}\nu^{2}}\frac{S_{\mathrm{point}}(\nu)}{\Omega}.$ (3)
Here, $c$ represents the speed of light, $k_{\rm B}$ is the Boltzmann constant, and $\Omega$ denotes the solid angle subtended by the telescope. The flux density $S_{\mathrm{point}}(\nu)$ of the background point source is modeled as a power-law spectrum with a spectral index $\eta$. This model is scaled to 150 MHz, represented as $S_{\rm point}(\nu)=S_{150}\left(\frac{\nu}{\nu_{150}}\right)^{\eta}$ Thyagarajan (2020), where $S_{150}$ represents the flux density at $150{\rm~MHz}$, with $\eta=-1.05$ Carilli et al. (2002).
[…]
$\delta T_{\mathrm{b}}(\hat{\bm{n}},z)\approx T_{0}(\hat{\bm{n}},z)\frac{1+\delta(z)}{T_{\rm K}(z)},$ (4)
where $T_{0}(\hat{\bm{n}},z)=-0.0085(1+z)^{1/2}T_{\gamma}(\hat{\bm{n}},z)$ is only related to the brightness temperature of the background radiation. Although $T_{\rm K}$ and $\delta$ actually describe the temperature and density fields in the direction of the background source, we will show in the subsequent analytical modeling that these physical quantities essentially originate from the isotropic halo mass function and the density and temperature profiles. Therefore, they are statistically independent of direction. Based on this, we omit the direction $\hat{\bm{n}}$ in our notation and simply write them as $T_{\rm K}(z)$ and $\delta(z)$.
@@
$P(\hat{\bm{n}},k_{\|},z)=T_{0}^{2}(\hat{\bm{n}},z)P_{21}(k_{\|},z),$ (5)
where $P_{21}(k_{\|},z)$ is the 1D power spectrum of the term $1+\delta(z)]/T_{\rm K}(z)$. It should be noted that $k_{\parallel}$ is, by definition, the Fourier variable of the comoving distance $r_{z}(z)$ along the line of sight. Therefore, it should theoretically vary with redshift $z$. However, in practical power spectrum analysis, due to the finite length of the observed or simulated region, the $k$ modes obtained from the discrete Fourier transform have statistical errors, especially in the high-$k$ region where large fluctuations are more likely to occur. To mitigate such uncertainties, it is common in power spectrum analysis to apply binning in $k$ space, where adjacent $k$ modes are grouped and averaged within the same bin. This procedure reduces statistical fluctuations in the estimated power spectrum, resulting in more reliable measurements. As a result, the binned $k_{\parallel}$ no longer corresponds to a specific redshift, but instead represents the average power over a fixed scale interval. Therefore, in subsequent analyses, we treat it as a redshift-independent variable to facilitate the unified comparison of power spectra and parameter constraints across different redshifts. We can obtain the 1D power spectrum by projecting the averaged three-dimensional (3D) power spectrum onto a specific direction,
$P_{21}(k_{\|},z)=P_{21,{\rm 1D}}(k,z)=\frac{1}{2\pi}\int_{k}^{\infty}k^{\prime}P_{21}(k^{\prime},z){\rm d}k^{\prime}.$ (6)
It should be noted that the 3D power spectrum $P_{21}(k,z)$ here is determined solely by the spatial distribution of neutral hydrogen, which depends on the density profile, temperature profile, and halo mass function of dark matter halos. In our model, these quantities are assumed to be isotropic, and thus $P_{21}(k,z)$ is also isotropic. In addition, redshift space distortions mainly originate from the peculiar velocities of the gas, which only cause a slight frequency drift in the absorption signal. We consider that the impact of this on the overall amplitude of the 1D power spectrum is relatively weak. Therefore, it is neglected in this work. Meanwhile, since we focus on scales smaller than 2 Mpc, the evolution of physical quantities along the line of sight with redshift can also be considered negligible on this scale. Based on these premises, we can project it along any direction to obtain the 1D power spectrum $P_{\rm 21,1D}(k,z)$, and the result along the line of sight is $P_{21}(k_{\parallel},z)$. Finally, through Eq. (5), we obtain the 1D power spectrum of the 21-cm forest $P(\hat{\bm{n}},k_{\|},z)$ along the line of sight. The information related to the direction of the background radio source including background source flux density $S_{150}$ is encoded in $T_{0}^{2}(\hat{\bm{n}},z)$.
@@ II.2 Halo model
where $P_{21}^{1h}(k,z)$ represents the contribution from a single halo and $P_{21}^{2h}(k,z)$ captures the contribution from two separate halos. The specific expressions for each term are given as follows:
@@
where $\rho_{21}(r,z,M)$ is the profile of $1+\delta(z)]/T_{\rm K}(z)$, including both the density profile and the temperature profile. In the following sections, we will expand on these components, detailing the specific forms of the density and temperature profiles.
@@ II.2.1 Halo mass function
Figure 1 illustrates the halo mass functions for CDM and WDM models at different redshifts. Since the 21-cm forest signal is mainly contributed by low-mass halos, we focus on halos with a minimum mass of $M_{\rm min}=10^{5}M_{\odot}$, corresponding to the Jeans mass scale for the redshifts of interest. The upper mass limit, $M_{4}$, is associated with a virial temperature of $T_{\rm vir}=10^{4}{\rm~K}$. Within this mass range, the WDM model predicts fewer halos than the CDM model.
Figure 1: Halo mass function in CDM model (solid lines) and WDM model (dashed lines) with $m_{\rm W}=6{\rm~keV}$ at different redshifts. The blue and green lines correspond to $z$ = 6 and 10, respectively.
@@ II.2.2 Density profile
Because of the gravitational influence, the gas density in the surroundings of halos is enhanced. Outside the virial radius, we assume the gas density profile aligns with that of the dark matter and can be determined using the infall model Barkana (2004). The gas density profiles in and around halos of different masses are plotted in the left panel of Fig. 2.
@@ II.2.3 Temperature profile
where $f_{\star}$ is the star formation efficiency, $f_{\rm coll}$ is the fraction of matter collapsed into atomic-cooling halos, and $f_{\rm X}$ is a normalization factor representing the uncertain x-ray productivity in the early Universe.
@@
The heating due to x-rays is dependent on the parameter $f_{\rm X}$, which characterizes the efficiency of x-ray production from early sources. For higher values of $f_{\rm X}$, the gas temperature in neutral regions is significantly raised, reducing the contrast between dense regions and the IGM in terms of their 21-cm absorption signatures. We show the temperature profiles for an unheated IGM with different masses in the right panel of Fig. 2. According to the recent Hydrogen Epoch of Reionization Array findings Abdurashidova et al. (2023), the temperature of IGM at redshift $z\sim$ 8 is constrained within the range $15.6{\rm~K}<T_{\rm K}<656.7{\rm~K}$, corresponding to $0.02<f_{\rm X}<0.6$. Thus for subsequent analysis, we adopt $f_{\rm X}=0.1$ as the fiducial model.
@@ III Forecasts
III.1 Radio-loud quasar population
With the advancement of optical and near-infrared detection technology, the frontier of quasar redshift has reached 7.64 Wang et al. (2021), while about $10\%$ of quasars exhibit strong radio emission Stern et al. (2000); Ivezic et al. (2002); Bañados et al. (2015); Gloudemans et al. (2021); Liu et al. (2021), making these radio-loud quasars ideal candidates for 21-cm forest observations. Using a physics-driven model Haiman et al. (2004); Niu et al. (2024), we can predict the future observations of radio-loud quasars. In combination with future infrared space telescopes, such as Euclid Scaramella et al. (2022) and the Nancy Grace Roman Space Telescope Spergel et al. (2015), we will be able to discover more high-redshift active galactic nuclei (AGNs). Additionally, by utilizing future powerful radio telescopes like the SKA, the New Extension in Nançay Upgrading Low-Frequency Array Munshi et al. (2024), and the upgraded Giant Metrewave Radio Telescope Gupta et al. (2017), we can filter out high-redshift radio-loud quasars from these AGNs.
We utilize the latest model of radio-loud quasars distribution presented in Ref. Niu et al. (2024) to simulate background point sources. The abundance of radio-loud quasars can be calculated as follows: